Can anyone help me with this one? I am have a hard time figuring this one out. I really need your help to pass this class. Thanks in advance.
True or false: The function "f(x) = 4x" grows four times faster than the function "g(x) = x". Explain.

Answer by RAY100(1637)

f(x) =4x,,,,g(x) =x
.
remember the slope intercept form,,y=mx +b
.
slope of f(x) =4/1,,,,,and slope of g(x) = 1/1
.
therefore for a move of one in x direction, f(x) grows to 4, but g(x) grows to 1
.
Hence the 4/1 growth ratio is true.